$ 150 weekly in saving account at 4.00% interest compounded saving any money.

a.how much will you have at the end of 5.5 years?

b. 250 monthly for the next 5 years discount 4%. what is the cost today?

c.15 % every year compounded quaterly, effective annual. what is the rate you are going to pay.
Answer #1

Using Microsoft Excel:
1.
=FV(4%/52,5.5*52,-150)=$47,964.41

2.
=PV(4%/12,12*5,-250)=$13,574.77

3.
=(1+quarterly annual rate/4)^4-1
=(1+15%/4)^4-1=15.87%